the cost of a gallon of gas increases from $1.46 to $3.53 over a period of 10 years. Use the formula r=(FP)1/n-1 to find the annual inflation rate r to the nearest tenth of a percent, where n is the number of years during which the value increases from P to F

Found 2 solutions by mananth, ikleyn:
Answer by mananth(16946)

Using the given formula, we can find the annual inflation rate:
r = (F/P)^(1/n) - 1
where P is the initial price ($1.46), F is the final price ($3.53), and n is the number of years (10).
Plug the values use calculator
r = (3.53/1.46)^(1/10) - 1
r = 1.140 - 1
r = 0.140
the annual inflation rate for the price of gas is approximately 14.0%.

Answer by ikleyn(52835)

.
the cost of a gallon of gas increases from $1.46 to $3.53 over a period of 10 years.
Use the formula r = (F/P)^(1/n)-1 to find the annual inflation rate r to the nearest
tenth of a percent,
where n is the number of years during which the value increases from P to F.
~~~~~~~~~~~~~~~~~~~


        Calculations in the post by @mananth are INCORRECT.
        I came to bring you a correct solution.


Using the given formula, we can find the annual inflation rate:

    r = (F/P)^(1/n) - 1

where P is the initial price ($1.46), F is the final price ($3.53), 
and n is the number of years (10).


Plug the values use calculator.

r = (3.53/1.46)^(1/10) - 1,

r = 0.092300632.


ANSWER. The annual inflation rate for the price of gas is approximately 9.2%, rounded as requested.

Solved (correctly).